# Anytime-valid simultaneous lower confidence bounds for the true discovery proportion

### v1

@@ 4.1 Non-decreasing anytime-valid simultaneous confidence bounds
Whether the proposed lower confidence bounds for the TDP are non-decreasing or not depends solely on the utilized e-process. While the probability of the e-process crossing a given threshold is bounded from above if the corresponding hypothesis is true, there are no guarantees that the e-process does not cross a given threshold multiple times if the hypothesis is false. That is, if $H_{I}$ is false, $\phi_{\alpha}^{[E,n]}$ might reject $H_{I}$ at a given time point $n$, but $\phi_{\alpha}^{[E,(n+k)]}$ might not reject (”accept”) $H_{I}$ at time point $n+k$, $k\in\mathbb{N}_{0}$. However, for the upper confidence bound to be non-decreasing, we require that only accept-to-reject decisions (ARD) are made, i.e., once an hypothesis is rejected by closed testing, it remains rejected.
@@
Second, ARD can be ensured by defining the local level-$\alpha$ test based on the maximum value of the so-far observed e-process, that is, $\phi_{\alpha}^{[E,n_{\ell}]}(H_{I})\coloneq\mathbb{I}(\max_{\ell\leq n}e_{I}^{[\ell]})$. While $\max_{\ell\leq n}E_{I}^{[\ell]}$ itself is not a valid e-process, $\phi_{\alpha}^{[E,n_{\ell}]}$ is a valid local level-$\alpha$ test due to Eq. (3). Alternatively, one could also define the local level-$\alpha$ test as $\phi_{\alpha}^{[E,\ddot{n}]}\coloneq\mathbb{I}(\ddot{e}_{I}^{[n]}\geq 1/\alpha)$, with
@@ 4.2 Choice of the e-process
where $\mathcal{L}(\cdot|\mu,\lambda)$ is the likelihood of the t-distribution with $\lambda$ degrees of freedom and non-centrality parameter $\mu$. Here, $n_{t}$ is the effective sample size used to compute the t-test statistic $t$. Furthermore, $\delta$ is a parameter that is chosen such that $lr_{n}(t)$ is an e-process. There are different ways to define $\delta$. Ly et al. (2024) recommend using a non-local moment prior on $\delta$, which is a two-bump prior with bumps at $-\delta_{min}$ and $\delta_{min}$, where $\delta_{min}$ is the minimal relevant effect size. They name the resulting e-process the ”mom” e-process. While the mom e-process does not necessarily have the largest growth rate in all scenarios, it’s growth-rate appears to be stable across different scenarios.
@@ 5 Simulations
$\boldsymbol{V}[a,b]$ =$=\begin{cases}1\qquad\text{if }a=b,\\ \rho\qquad\text{else.}\end{cases}$
@@
We have considered nine different discovery sets with TDP of $\pi_{1}\in\{0.1,0.2,\ldots,0.9\}$. For each discovery set we have computed the $1-\alpha$ anytime-valid simultaneous lower confidence bounds for the TDP, $\hat{\pi}_{1}$, based on the three processes, with $\alpha=0.2$. The bounds are based on at least 11 observations. The results are based on 1000 Monte-Carlo iterations.
@@
Figure 1: Proportion of iterations with $\hat{\pi}_{1}>\pi_{1}$ for $\rho=0.2$ over time, where $\hat{\pi}_{1}$ is the anytime-valid simultaneous lower confidence bound for the true discovery proportion $\pi_{1}$. The three different utilized local level-$\alpha$ e-tests are based on the mom e-process values (mom), the mom e-process values adapted to only allow accept-to-reject decisions (mom_ARD) and the e-to-p process values based on the mom e-process (e-to-p), respectively. The black solid line corresponds to the significance level $\alpha=0.2$. Results are based on $1000$ iterations.
@@
Figure 2: Average of the anytime-valid simultaneous lower confidence bound $\hat{\pi}_{1}$ for the true discovery proportion $\pi_{1}$ over time. The three different utilized local level-$\alpha$ e-tests are based on the mom e-process values (mom), the mom e-process values adapted to only allow accept-to-reject decisions (mom_ARD) and the e-to-p process values based on the mom e-process (e-to-p), respectively. The black solid line corresponds to $\pi_{1}$. Results are based on $1000$ iterations.
@@ 6 Discussion
The presented procedure has been derived under the assumption that the utilized e-processes are defined with respect to the natural filtration. While this seems to be a sensible requirement in practical settings, it is not a necessary requirement. Indeed, the procedure can be applied using e-processes for any other filtration $(\mathcal{G}_{n})_{n\geq 0}$ as well. In this case, the procedure only returns valid results for $\mathcal{G}$-stopping times. All e-processes utilized in the procedure need to be e-processes with respect to the same filtration.

### v2

@@ 4.1 Non-decreasing anytime-valid simultaneous confidence bounds
Whether the proposed lower confidence bounds for the TDP are non-decreasing or not depends solely on the utilized e-process. While the probability of the e-process crossing a given threshold is bounded from above if the corresponding hypothesis is true, there are no guarantees that the e-process does not cross a given threshold multiple times if the hypothesis is false. That is, if $H_{I}$ is false, $\phi_{\alpha}^{[E,n]}$ might reject $H_{I}$ at a given time point $n$, but $\phi_{\alpha}^{[E,(n+k)]}$ might not reject (”accept”) $H_{I}$ at time point $n+k$, $k\in\mathbb{N}_{0}$. However, for the upper confidence bound to be non-decreasing, we require that only accept-to-reject decisions (ARD) are made, i.e., once an hypothesis is rejected, it remains rejected.
@@
Second, ARD can be ensured by defining the local level-$\alpha$ test based on the maximum value of the so-far observed e-process, that is, $\phi_{\alpha}^{[E,n_{\ell}]}(H_{I})\coloneq\mathbb{I}(\max_{\ell\leq n}e_{I}^{[\ell]}\geq 1/\alpha)$. While $\max_{\ell\leq n}E_{I}^{[\ell]}$ itself is not a valid e-process, $\phi_{\alpha}^{[E,n_{\ell}]}$ is a valid local level-$\alpha$ test due to Eq. (3). Alternatively, one could also define the local level-$\alpha$ test as $\phi_{\alpha}^{[E,\ddot{n}]}\coloneq\mathbb{I}(\ddot{e}_{I}^{[n]}\geq 1/\alpha)$, with
@@ 4.2 Choice of the e-process
where $\mathcal{L}(\cdot|\mu,\lambda)$ is the likelihood of the t-distribution with $\lambda$ degrees of freedom and non-centrality parameter $\mu$. Here, $n_{t}$ is the effective sample size used to compute the t-test statistic $t$. Furthermore, $\delta$ is a parameter that is chosen such that $lr_{n}(t)$ is an e-process. There are different ways to define $\delta$. Ly et al. (2024) recommend using a non-local moment prior on $\delta$, which is a two-bump prior with bumps at $\delta_{min}$ and $\delta_{min}$, where $\delta_{min}$ is the minimal relevant effect size. They name the resulting e-process the ”mom” e-process. While the mom e-process does not necessarily have the largest growth rate in all scenarios, it’s growth-rate appears to be stable across different scenarios.
@@ 5 Simulations
$\boldsymbol{V}[a,b]$ $=\begin{cases}1\qquad\text{if }a=b,\\ \rho\qquad\text{else.}\end{cases}$
@@
We have considered nine different rejection sets with TDP of $\pi_{1}\in\{0.1,0.2,\ldots,0.9\}$. For each rejection set we have computed the $1-\alpha$ anytime-valid simultaneous lower confidence bounds for the TDP, $\hat{\pi}_{1}$, based on the three processes, with $\alpha=0.2$. The bounds are based on at least 11 observations. The results are based on 1000 Monte-Carlo iterations.
@@
Figure 1: Empirical non-coverage rate over time for $\rho=0.2$. The three different utilized local level-$\alpha$ e-tests are based on the mom e-process values (mom), the mom e-process values adapted to only allow accept-to-reject decisions (mom_ARD) and the e-to-p process values based on the mom e-process (e-to-p), respectively. The black solid line corresponds to the significance level $\alpha=0.2$. Results are based on 10001000 iterations.
@@
Figure 2: Average of the anytime-valid simultaneous lower confidence bound $\hat{\pi}_{1}$ for the true discovery proportion $\pi_{1}$ over time for $\rho=0.2$. The three different utilized local level-$\alpha$ e-tests are based on the mom e-process values (mom), the mom e-process values adapted to only allow accept-to-reject decisions (mom_ARD) and the e-to-p process values based on the mom e-process (e-to-p), respectively. The black solid line corresponds to $\pi_{1}$. Results are based on 10001000 iterations.
@@ 6 Discussion
The presented procedure has been derived under the assumption that the utilized e-processes are defined with respect to the natural filtration. While this seems to be a sensible requirement in practical settings, it is not a necessary requirement. Indeed, the procedure can be applied using e-processes for any other filtration $(\mathcal{G}_{n})_{n\geq 0}$ as well. In this case, the procedure only returns valid results for $\mathcal{G}$-stopping times.